OpenStudy (anonymous):

Let f(x)=2x+3 and g(x)-x^2-x find the value of (f+g)(1/2)

OpenStudy (anonymous):

is it 15/4 not sure *mental calculation

OpenStudy (cwrw238):

add them and plug in x = 1/2

OpenStudy (anonymous):

You have f(x) as well of g(x). You just add them together like you'd add anything. f(x)+g(x)=(2x+3)+(x^2-x)=x^2+x+3 then plug in x=1/2 for the x values.

OpenStudy (anonymous):

as well as*

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

i pluged in the 1/2 for the x's but i dont get it

OpenStudy (anonymous):

what dont you get just put 1/2 and add up both

OpenStudy (anonymous):

 Help at all?